OpenStudy (anonymous):

Would the answer to h(x) = x^2 - 5 find h(-3) be -14?

OpenStudy (anonymous):

(-3)^2-5= 9-5= 4

OpenStudy (anonymous):

when you square a negative it becomes positive (-3)*(-3)=9

OpenStudy (anonymous):

Oh I forgot to do that, thank you.